# A Comprehensive Overview of Large Language Models

### 27 Dec 2023 v7

@@ VII Applications
The application of Large Language Models (LLMs) has become a hot topic in both AI-related research communities and industries, with many emerging uses being discovered and explored daily. These models, capable of understanding and generating human-like text, have found meaningful applications across a variety of fields. This section will provide an overview of LLMs’ applications in medicine, education, science, mathematics, law, finance, robotics, and coding. While each of these domains poses different challenges, LLMs have opportunities to make significant contributions given their wide-ranging scope of applicability.
@@
General Purpose: One of the most significant applications of LLMs is as a general-purpose tool for a wide variety of tasks, even those that have not been specifically trained for [433]. This is due to their inherent ability to understand, generate, and manipulate human-like text in a contextually relevant manner. This allows them to perform tasks ranging from simple language translation and question-answering to more complex tasks like summarization, text generation, and even programming help [434]. The utility of LLMs is further enhanced by their ability to adapt to the specific style and tone of the text they are processing, making the outputs more user-friendly and context-aware. In everyday applications, LLMs can be used as personal assistants, helping users draft emails or schedule appointments [435]; they can also be deployed in customer service to handle common questions, freeing up human resources for more complex issues; or applied to generate content for digital platforms like websites, by creating human-like text based on given prompts [436]. Moreover, LLMs play a crucial role in data analysis, where they can filter large volumes of text data, summarize key points, and find patterns that would take humans much longer time to identify [437]. Despite their wide-ranging applications, it is essential to remember that LLMs, like any AI system, are only as good as the data they have been trained on. We should always use LLMs carefully since they can unintentionally reproduce biases in their training data, leading to potentially unfair or inaccurate results.
@@
Others: The application of LLM in coding is introduced in Section III-A2. For the utility of LLMs in robotics, please refer to Section III-D.
VIII Summary and Discussion
VIII-A Architecture
Layer Normalization is found to have a significant effect on the performance and training stability of LLMs. Pre-norm, that is normalizing inputs rather than outputs, is more common among LLMs stabilizing the training [6, 126, 104]. BLOOM [13] and AlexaTM [122] utilize an additional layer normalization before embedding layer to stabilize the training of large-scale models, while the model’s zero-shot generalization ability can be negatively impacted [13]. However, another study [33] finds that pre-norm degrades fine-tuned model performance as compared to post-norm, and there are no stability benefits of pre-norm beyond the 100B scale. Therefore, GLM-130B [33] used deep-norm which is a variant of post-norm for better downstream task performance after fine-tuning.
Positional Encoding effect performance and training stability of LLMs like other building blocks of a model. BLOOM [13] finds ALiBi outperforming learned and rotary positional encodings. Contrary to this, GLM-130B [33] identifies rotary positional encoding better than ALiBi. So, there is no conclusion in literature about the positional encodings yet.
Parallel Attention where attention and feed-forward layers are parallel to each other rather than sequential in transformer block has shown to reduce training time by 15%. There is no evidence of performance drop due to this change in literature and used by the models PaLM [15], GPT-NeoX [115], and CodeGen [130].
Mixture of Experts allows easily scaling model to trillion of parameters [129, 118]. Only a few experts are activated during the computation making them compute-efficient. The performance of MoE models is better than the dense models for the same amount of data and requires less computation during fine-tuning to achieve performance similar to the dense models as discussed in [118]. MoE architectures are less prone to catastrophic forgetting, therefore are more suited for continual learning [129]. Extracting smaller sub-models for downstream tasks is possible without losing any performance, making MoE architecture hardware-friendly [129].
Sparse vs Dense Activated GPT-3 [6] uses sparse transformers [63] whereas GLaM [118] and PanGu-$\sum$ [129] use MoE [119] architecture to lower computational costs and increase the model size and capacity. According to the literature, sparse modules do not degrade the model’s performance [63]. However, more experiments are required to verify this statement.
VIII-B Training Strategies
Training models at a huge scale require some tricks to reduce training costs, avoid loss divergence and achieve better performance. We summarize and discuss some of these key tricks used in different LLMs.
Mixed Precision is a famous method for LLMs to reduce memory usage and improve training efficiency. In mixed precision, forward and backward passes are performed in FP16 format whereas optimizer states and master weights are kept in FP32 format [474]. A drawback associated with this format change is training instability due to a smaller value range resulting in loss spikes [33]. An alternative to FP16 is BF16 which has a comparatively larger range and performs some precision-sensitive operations like gradient accumulation and softmax in FP32 [13]. BF16 has better performance and training stability but uses more memory and is supported on specific hardware, for example, A100 GPUs. Therefore, its adoption in LLMs is limited.
Training Instability is a common issue in LLMs where loss divergence or spiking is observed multiple times during training. This happens in the presence of gradient clipping [15]. To mitigate this problem, many approaches suggest restarting training from an earlier checkpoint [15, 33, 118], skipping 200-500 earlier data batches at the point of divergence in [15] and re-shuffling batches in [118]. The embedding layer gradient shrink proves to further stabilize the training as its gradient norm is significantly larger than the other layers [33]. Another suggestion to improve training stability for larger models is not to use biases in dense and norm layers as in [15].
Weight Initialization plays a significant role in model convergence and training stability. GPT-NeoX [115] initializes feed-forward layers before residuals with $\frac{2}{L\sqrt{d}}$ as in [144] and other layers with small initialization scheme [475]. This avoids activations growing exponentially with the increasing depth. MT-NLG [114] found higher variance for weight initialization leads to unstable training, hence validating small initialization scheme [475]. Various models perform random weight initialization which can cause bad initialization, Galactica [138] suggests a longer warmup to negate the effect.
Learning Rate is important for stable training. It is suggested to use a lower value [13, 15, 124] with warmup and decay (cosine or linear). Usually, the learning rate is within the range $1e^{-4}$ to $8e^{-4}$. Moreover, MT-NLG (530B) [114] and GPT-NeoX (20B) [115] suggest interpolating learning rates based on the model size using the GPT-3 [6] models ranging between 13B and 175B. This avoids tuning the learning rate hyperparameter.
Training Parallelism 3D parallelism, a combination of data, pipeline and tensor parallelism, is the most utilized training parallelism approach in LLMs [33, 15, 14, 13, 114, 112, 109]. In addition to the 3D parallelism, BLOOM [13] uses zero optimizer [37] to shard optimizer states. PanGu-$\alpha$ [104] and PanGu-$\Sigma$ [129] go beyond the 3D parallelism and apply 5D parallelism which additionally contains optimizer parallelism and rematerialization.
Mode Switching adds task-related tokens at the beginning of the text during training. These tokens refer to the natural language understanding and natural language generation tasks which are shown to improve the downstream task performance in [89, 124, 122]. During fine-tuning and inference, tokens are appended based on the downstream tasks.
Controllable Text Generation Generating credible and controlled text from a pre-trained model is challenging. GPT-3 [6] and other LLMs use in-context learning to control generated text. While in-context learning helps in controlling the generated text, ERNIE 3.0 Titan [35] suggests using adversarial loss to rank its generated text for credibility and soft prompts such as genre, topic, keywords, sentiment, and length for better control on generated text.
VIII-C Pre-Training vs Instruction Tuning
While pre-training is important for the generalization of LLMs, instruction-tuning improves the performance of LLMs further and makes them useable. Therefore, it is suggested to perform instruction fine-tuning of pre-trained LLMs to use them effectively [16, 18, 20, 93, 157].
VIII-D Supervised Models vs Generalized Models
VIII-E Zero-Shot vs Few-Shot
VIII-F Encoder vs Decoder vs Encoder-Decoder
IX Challenges and Future Directions
@@ X Conclusion
XI Versioning
We keep track of the versions of this paper we release as the content updates.
Version 1.0: We covered 30 pre-trained models and 6 instruction-tuned models, including their overview, findings, training, and evaluation datasets, and discussed important architectural and training tricks by various LLMs.
Version 2.0: Further pre-trained LLMs added along with discussion on on self-instruct LLMs. Categorized LLMs according to the application, provided descriptions of widely used evaluation datasets, added a section on robotics, and extended discussion in section VIII. Tables have been updated.
Version 3.0: Added sections on Alignment tuning and multimodal LLMs. A performance comparison table on various benchmarks and datasets. Added LLaMA-2 and PaLM-2.
Version 4.0: Tables on training and evaluation datasets, a sub-section on increasing context window, and minor improvements.
Version 5.0: Added sections on augmented LLMs and challenges and future directions.
Version 6.0: Minor improvements in abstract, introduction, and conclusion.
Version 7.0: Sections on efficient LLMs and applications.
Note: If you find any mistakes, or have issues and conflicts with the writing in this paper, please email us. We welcome suggestions to improve this paper.

### 20 Feb 2024 v8

@@ VI Applications
Applying Large Language Models (LLMs) to a variety of downstream tasks has become a popular trend in both AI-related research communities and industries, with many emerging uses being discovered and explored daily. LLMs, which are capable of understanding and generating human-like text, have found meaningful applications across a variety of fields. This section provides an overview of LLM applications in medicine, education, science, mathematics, law, finance, robotics, and coding. While each of these domains pose different challenges, LLMs open up opportunities to make significant contributions to these domains through their generalizability.
@@
General Purpose: LLMs are being widely considered as general-purpose tools for a wide variety of tasks [420]. This is due to their inherent ability to understand, generate, and manipulate human-like text in a contextually relevant manner. This allows them to perform tasks ranging from simple language translation and question-answering to more complex tasks like summarization, text generation, and even programming help [421]. The utility of LLMs is further enhanced by their ability to adapt to the specific style and tone of the text they are processing, making the outputs more user-friendly and context-aware. In everyday applications, LLMs can be used as personal assistants, helping users draft emails or schedule appointments [422]; they can also be deployed in customer service to handle common questions; or applied to generate content for digital platforms like websites, by creating human-like text based on given prompts [423]. Moreover, LLMs play a crucial role in data analysis, where they can filter large volumes of text data, summarize key points, and find patterns that would take humans much longer to identify [424]. Despite their wide-ranging applications, it is essential to remember that LLMs, similar to any AI system, are only as good as the data they have been trained on.
@@
Robotics: In robotics research, LLMs have promising applications, such as enhancing human-robot interaction [28, 461, 462, 463], task planning [226], motion planning [235], navigation [235, 464], object manipulation [225], personalized robots [465], etc. LLMs enable robots to understand the environment effectively and generate plans to complete tasks collaboratively [229, 26]. They can facilitate continuous learning by allowing robots to access and integrate information from a wide range of sources, helping robots acquire new skills, adapt to changes, and refine their paths [213, 222, 223].
VII Challenges and Future Directions